laylasnii:

Help Please Determine the answer to 3 + (−5) and explain the steps using a number line.

Timmyspu:

Do you have a screenshot of the number line you have to use?

Joe348:

\[3+(5)=-2\]

laylasnii:

@timmyspu wrote:
Do you have a screenshot of the number line you have to use?
there is no number line just a blank box i have the answer i just need help explaining how i got it

Joe348:



Joe348:

A number line like this?

laylasnii:

@joe348 wrote:
A number line like this?
i have no idea it shows no number line

laylasnii:

laylasnii:

thats what it looks like

Timmyspu:

@joe348 wrote:
\[3+(5)=-2\]
3+(-5)=-2 Sorry today is just not my day.

laylasnii:

I need help explaining how i got it

laylasnii:

it's confusing sorry guys

Joe348:

you could say take out the + sign and just do \[3-5\]

laylasnii:

@joe348 wrote:
you could say take out the + sign and just do \[3-5\]
tysm

laylasnii:

this is what i put

Timmyspu:

Wait is this a test or a quiz?

laylasnii:

quiz

Joe348:

oh waitttt

Joe348:

dont submit your answer as that though

laylasnii:

@joe348 wrote:
dont submit your answer as that though
i didn't

Joe348:

just say this

Timmyspu:

Oh then Joe we can not help out because qc rules say we are unable to help with any test or quiz assignment. I am sorry. For next time please do not post a question related to a test or quiz. Thank you!

Joe348:

\[3+(-5)= 3-5= -2\]

Joe348:

Either way she has the answer now but yeah, we can not help with quiz or tests questioncove.com/rules

laylasnii:

Well Thank you guys so much for helping and i will make sure not to use any question from a test or quiz

Timmyspu:

https://questioncove.com/rules

Joe348:

omg ty timmy😂

Timmyspu:

Your welcome! I saw you were having a bit of trouble xd. Alright just please do not post another question related to a quiz or test.

laylasnii:

will do again thx

Timmyspu:

Any time!